Tier1 Operations Report 2018-04-30

From GridPP Wiki
Jump to: navigation, search

Editing Tier1 Operations Report 2018-04-30 Bold textItalic textInternal linkExternal link (remember http:// prefix)Level 2 headlineEmbedded fileFile linkIgnore wiki formattingYour signature with timestampHorizontal line (use sparingly)

RAL Tier1 Operations Report for 30t April 2018

Review of Issues during the week 23rd April to the 30th April 2018.
  • The proposed emergency chassis intervention was brought forward (23/5/18 - 13:00), when the chassis began to intermittently reboot itself. This intervention was completed successfully and without further incident.
  • The scheduled firewall replacement (25/4/18), appeared to be completed successfully. However, during the afternoon the connection between the Tier1 network and the RAL core was failing intermittently. The problem seemed to get worse as the afternoon progressed. Due to disruptions to connectivity both on and off site, caused (we think) by running on the new secondary firewall, we flipped the Master-Slave router pair so that the Master was running on the standby side. The change occurred at approximately 17:00 on the Wednesday and normal service was subsequently restored. There is currently an ongoing investigation as to why the firewalls did not fail-over correctly. There is an "at risk" scheduled outage for the 2nd May 2018 when Networking/DS will be performing another firewall switch-over to ensure that this issue has been resolved.
Current operational status and issues
  • None
Resolved Castor Disk Server Issues
  • None.
Ongoing Castor Disk Server Issues
  • gdss737 (lhcbDst- D1T0) - Currently out of production awaiting investigation by the CASTOR and Fabric teams.
Limits on concurrent batch system jobs.
  • CMS Multicore 550
Notable Changes made since the last meeting.
  • None.
Entries in GOC DB starting since the last report.
  • None
Declared in the GOC DB
Downtime id Hosts Start Time End Time Severity Description
25263 All RAL 02/05/2018 07:30 02/05/2018 08:30 SCHEDULED WARNING Firewall fail-over test. Within this time window there is the expectation of a break in connectivity while connections are moved between the two firewalls.
Advanced warning for other interventions
The following items are being discussed and are still to be formally scheduled and announced.

Listing by category:

  • Echo
    • Applying minor CEPH update to fix the "backfill" bug. Increase number of placement groups and add more capacity.
  • Castor:
    • Update systems to use SL7 and configured by Quattor/Aquilon. (Tape servers done)
    • Move to generic Castor headnodes.
  • Networking
    • Extend the number of services on the production network with IPv6 dual stack. (Done for Perfsonar, FTS3, all squids and the CVMFS Stratum-1 servers).
    • Replacement (upgrade) of RAL firewall.
  • Internal
    • DNS servers will be rolled out within the Tier1 network.
  • Infrastructure
    • Testing of power distribution boards in the R89 machine room is being scheduled for some time late July / Early August. The effect of this on our services is being discussed.
Open GGUS Tickets (Snapshot during morning of the report)
Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
117683 none on hold less urgent 18/11/2015 28/03/2018 Information System CASTOR at RAL not publishing GLUE 2 EGI
124876 ops on hold less urgent 07/11/2016 13/11/2017 Operations [Rod Dashboard] Issue detected : hr.srce.GridFTP-Transfer-ops@gridftp.echo.stfc.ac.uk EGI
127597 cms on hold urgent 07/04/2017 16/04/2018 File Transfer Check networking and xrootd RAL-CERN performance EGI
133992 atlas in progress less urgent 12/03/2018 19/04/2018 File Transfer RAL-LCG2-ECHO: No such file or directory EGI
134468 cms waiting for reply top priority 09/04/2018 24/04/2018 CMS_AAA WAN Access Xrootd redirector not seeing some files in ECHO WLCG
134494 atlas waiting for reply urgent 11/04/2018 29/04/2018 Storage Systems json space reporting not updated WLCG
134619 cms in progress urgent 19/04/2018 26/04/2018 CMS_SAM tests Problems reading data from ECHO WLCG
134685 dteam in progress less urgent 23/04/2018 23/04/2018 Middleware please upgrade perfsonar host(s) at RAL-LCG2 to CentOS7 EGI
134769 cms in progress urgent 26/04/2018 27/04/2018 CMS_Data Transfers Transfers from RAL_Disk to Florida are failing WLCG
GGUS Tickets Closed Last week
Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
132589 lhcb verified very urgent 21/12/2017 23/04/2018 Local Batch System Killed pilots at RAL WLCG
134136 atlas closed less urgent 20/03/2018 25/04/2018 File Transfer RAL-LCG2: [SE][Ls][SRM_INVALID_PATH] No such file or directory EGI
134342 cms solved urgent 29/03/2018 26/04/2018 CMS_Data Transfers CERN->MSS rate test WLCG
134413 none closed less urgent 05/04/2018 24/04/2018 VO Specific Software /cvmfs/larsoft.opensciencegrid.org not mounting EGI
134438 snoplus.snolab.ca closed less urgent 06/04/2018 26/04/2018 File Transfer FTS - transfer rate success drop EGI
134488 snoplus.snolab.ca solved urgent 10/04/2018 25/04/2018 File Transfer FTS failures to RAL EGI
134599 t2k.org solved less urgent 18/04/2018 24/04/2018 File Access Unable to reach 25 files at RAL which are ONLINE AND NEARLINE EGI
134627 cms solved urgent 19/04/2018 24/04/2018 CMS_Data Transfers Transfer backlogs WLCG
134703 cms solved urgent 23/04/2018 26/04/2018 CMS_Data Transfers Transfer failing from RAL_Disk WLCG
134731 lhcb solved very urgent 25/04/2018 26/04/2018 File Access could not open connection to srm-lhcb.gridpp.rl.ac.uk:8443 WLCG
134737 cms solved urgent 25/04/2018 30/04/2018 CMS_SAM tests SAM CE test failing at T1_UK_RAL WLCG
134744 cms solved top priority 25/04/2018 26/04/2018 File Transfer Zero Phedex Transfers - via RAL FTS service on certain links EGI
Availability Report
Target Availability for each site is 97.0% Red <90% Orange <97%
2018-04-22 100 100 100 100 100 100
2018-04-23 100 100 100 100 100 100
2018-04-24 100 100 100 99 100 100
2018-04-25 82 85 100 87 100 79.5053
2018-04-26 100 99 99 96 100 100
2018-04-27 100 98 98 96 100 100
2018-04-28 98 100 100 100 100 100
2018-04-29 100 100 100 100 100 100
Hammercloud Test Report
Target Availability for each site is 97.0% Red <90% Orange <97%

Key: Atlas HC = Atlas HammerCloud (Queue RAL-LCG2_UCORE, Template 841); CMS HC = CMS HammerCloud

Day Atlas HC CMS HC Comment
2018/04/16 100 100
2018/04/17 100 99
2018/04/18 97 95
2018/04/19 82 99
2018/04/20 100 100
2018/04/21 100 99
2018/04/22 100 99
Notes from Meeting.
  • None yet


This is a minor edit  Watch this page

Please note that all contributions to GridPP Wiki may be edited, altered, or removed by other contributors. If you do not want your writing to be edited mercilessly, then do not submit it here. You are also promising us that you wrote this yourself, or copied it from a public domain or similar free resource (see GridPP Wiki:Copyrights for details). Do not submit copyrighted work without permission!

  Cancel | Editing help (opens in new window)

Navigation menu Darren Moore 213b9e8222TalkPreferencesWatchlistContributionsPageDiscussionReadEditView historyWatch

Search Search GridPP Homepage Meetings News Navigation Main page Recent changes Random page Tools Privacy policyAbout GridPP WikiDisclaimersPowered by MediaWiki